OpenStudy (anonymous):

Please help someone?? Andrew is financing a condominium for $543,000. He gets a 30-year, fixed rate mortgage with a rate of 5.475%. How much interest will he pay over the life of his loan? Round to the nearest dollar.

jimthompson5910 (jim_thompson5910):

use this calculator http://www.bankrate.com/calculators/mortgages/loan-calculator.aspx to figure out how much the monthly payment is. Tell me what you get.

OpenStudy (anonymous):

I got 3,074.58 but it isn't one of my answer choices..

jimthompson5910 (jim_thompson5910):

I know, but that will help you get there

jimthompson5910 (jim_thompson5910):

so you make monthly payments of $3,074.58 for 360 months (note: 360 months = 30 years) this means you make a total of (3074.58)*(360) = 1,106,848.80 dollars in payments over the entire course of the mortgage now you subtract off the principal to find the interest Total Interest = (Total Amount Paid) - (Total Amount Loaned Out) = (1106848.80) - (543000) = 563848.80 so, Total Interest = $563,848.80

jimthompson5910 (jim_thompson5910):

oh and you round to the nearest dollar to get Total Interest = $563,849

OpenStudy (anonymous):

Your awesome and you explained it very well thank you!

jimthompson5910 (jim_thompson5910):

you're welcome, glad it helped

OpenStudy (anonymous):

Can you help with a few more?